# Barycenter kinematics in Local Group analogues

### v2

@@ Abstract
We present a new way to identify systems similar to our Local Group (LG) of galaxies in cosmological simulations. Our method uses as a new constraint the speed and direction of the LG’s center of mass, which we can measure accurately from cosmic microwave background data. When we apply these criteria to different cosmological simulations and compare the results with traditional selection methods, we find statistically significant differences. Our approach produces simulated galaxy pairs where the relative M31 velocity is less radial ($2\%$ to $10\%$ difference over the mean) and more tangential ($1\%$ to $4\%$ difference over the mean) than in cases that do not take into account the barycenter speed. The radial change pattern appears consistently across all cosmological models we test, showing that matching the observed barycenter velocity has a measurable effect when modeling and interpreting Local Group-like systems.
@@ 1 Introduction
We identified an important measurement that previous studies haven’t used when defining LG-like systems. This measurement is the LG’s barycenter velocity ($\mathbf{v}_{b}$), which we can determine precisely from cosmic microwave background (CMB) data (Planck Collaboration et al., 2020). We use this constraint in two ways. First, we consider its magnitude ($v_{b}$). Second, we examine the direction, measured as the angle ($\theta$) between the barycenter velocity vector and the position vector from MW to M31 ($\mathbf{r}$), as shown in Figure 1. We express this angle more conveniently as $\mu=\cos\theta$, which comes from the dot product of the unit vectors for $\mathbf{v}_{b}$ and $\mathbf{r}$.
@@
Our paper is organized as follows: Section 2 covers the observational constraints, Section 3 describes our simulations and the different cosmological models. In Section 4, we explain how we select LG analogues using both traditional criteria and our new approach. Section 5 presents and discusses our findings, and Section 6 summarizes our conclusions.
Figure 1: Diagram showing the alignment between key quantities used to select LG analogues. $\mathbf{v}_{b}$ represents the barycenter velocity of the LG and $\mathbf{r}$ shows the relative position from M31 to the MW
@@
$v_{r,\text{M31}}$ $-109.3\pm 4.4$ $\text{kms}^{-1}$ van der Marel et al. (2012)
@@
$\mu$ $-0.88\pm 0.01$ adim. This work
@@ 2 Observational constraints
Figure 2: Left panel: cumulative distribution of $\mu$, the cosine of the angle between the barycenter velocity of the LG and the relative position vector from M31 to the MW. Right panel: Cumulative distribution for the barycenter speed $v_{b}$ of the Local Group.
@@ 3 Simulation
We begin with the primary cosmology (c000), which uses the Planck2018 $\Lambda$CDM model with massive neutrinos. We then analyzed 91 additional cosmologies. These include: c009 (without neutrino mass), c019 and c020 (with varied neutrino species counts), and four secondary cosmologies (c001-c004) that test different scenarios such as low $\Omega_{c}$, wCDM (c002), high $N_{\text{eff}}$, and low $\sigma_{8}$.
@@ 4 Selection of LG analogues
We use halos identified with the CompaSO Halo Finder (Hadzhiyska et al., 2021), with selection criteria similar to those in Forero-Romero & Sierra-Porta (2022). We selected halos with maximum circular velocities $V_{\text{max}}>200$ km s-1.
From these, we identify isolated pairs as two halos, A and B, that are each other’s nearest neighbors. We also require that no third halo with a maximum circular velocity greater than halo B’s can be found within three times the pair’s separation distance. We define halo A as the one with the greater maximum circular velocity, which we designated as M31. This forms our main sample, typically containing about $8\times 10^{5}$ pairs per simulation box.
@@
  • Total pair mass between $[0.5,5.0]\times 10^{12}$ M
@@
  • Barycenter speed greater than the observed value ($v_{b}>620$ km s-1)
  • Alignment value less than the observed value ($\mu<-0.88$)
@@
The traditional subsample typically contains about $2\times 10^{5}$ pairs per simulation box, while the realistic sample contains about $10^{4}$ pairs.
Figure 3: Impact of our proposed barycenter velocity constraints on the mean velocity distributions of M31. Only cosmologies with statistically significant differences are shown. Dashed lines and shaded regions represent the average and standard deviation of the mean values across traditional pair distributions for all cosmologies. The secondary y-axis shows the percentage difference in mean values compared to the average mean of traditional pairs.
Figure 4: Statistical significance of the differences between realistic and traditional selection criteria for radial velocity $v_{r}$ (left column) and tangential velocity $v_{t}$ (right column). The $p$-values shown are from two-sample Kolmogorov-Smirnov tests across all cosmologies.
Figure 5: Cumulative distributions of M31 velocities relative to the MW. The plots compare observed values with traditional and realistic LG analogues for the most statistically significant cosmologies. Left panel shows radial velocity ($v_{r}$) distributions, and right panel shows tangential velocity ($v_{t}$) distributions.
@@ 5.1 Base cosmology
Figure 2 shows the cumulative distributions of $\mu$ and $v_{b}$ in the base cosmology (c000). These distributions illustrate the key features of our barycenter velocity constraints used to select LG analogues. The $\mu$ distribution reveals that $\mathbf{v}_{b}$ and $\mathbf{r}$ in simulated LG pairs tend to be either aligned or anti-aligned. We confirm this by comparing against a uniform distribution using a KS test, which shows statistically significant differences ($p<0.05$). Using the same statistical method, we also find that the distribution is not symmetric with a slight tendency towards anti-alignment.
@@ 5.2 All cosmologies
Our barycenter velocity constraints consistently affect the distributions of both $v_{t}$ and $v_{r}$ compared to traditional constraints across different cosmological models. Figure 3 shows how the mean values of tangential and radial velocities change across all cosmologies with statistically significant differences according to KS tests. Among the 93 cosmologies we study, the differences in radial velocity distribution are significant in 86 models, while tangential velocity distribution shows significant differences in 11 cosmologies. The effect on total mass ($M_{\text{LG}}$) is also significant in 11 cosmologies. Table 2 summarizes these results.
Table 2: Average of mean values across cosmologies for different properties. Only cosmologies with statistically significant differences are included in these averages. The columns show, respectively: the physical observable, the sample type, the average value across cosmologies, the percentage difference between the mean of the samples, and the number of cosmologies (out of 93) where the difference was statistically significant.
@@
Observable Sample Mean Mean difference (%) Significant cosmologies
$v_{r}$ / km s-1 Traditional $-85.39$
Realistic $-81.05$ $-4.86$ 86/93
$v_{t}$ / km s-1 Traditional $140.11$
Realistic $141.93$ $1.42$ 11/93
$M_{\text{LG}}$/ $10^{12}\text{M}_{\odot}$ Traditional $6.41$
Realistic $6.27$ $-0.88$ 11/93
We next investigate which cosmological parameters most strongly determine whether the difference in $v_{t}$ and $v_{r}$ between traditional and realistic pairs is significant. We use a Random Forest algorithm with the KS test p-values as the target variable. We performed a k-fold stratified cross validation and obtained a mean weighted average F1 score of 0.88 for $v_{r}$ and 0.79 for $v_{t}$. It is important to note that due to the class imbalance, the model is skewed towards correctly classifying the majority class, that is, the significant cosmologies in $v_{r}$ and the not significant in $v_{t}$. Figure 6 shows the results, with $\sigma_{8\text{cb}}$ (that is $\sigma_{8}$ taking only into account dark matter and baryons) emerging as the most influential parameter for $v_{r}$, and $\Omega_{\text{cdm}}$ (the dark matter density parameter) for $v_{t}$. Figure 7 shows the relationships between the most influential parameters for $v_{r}$ and $v_{t}$. Although we identify this correlation, a deeper exploration of the physical mechanisms explaining why these parameters most strongly affect each velocity component differences remains for future work.
@@
Figure 6: Relative importance of various cosmological parameters for predicting whether the difference in M31’s velocities between traditional and realistic pairs is statistically significant. The top panel shows the parameter importance for $v_{r}$, and the bottom panel for $v_{t}$. Importance is calculated using a Random Forest algorithm.
@@
Figure 7: Scatter matrix showing pairwise relationships between the most important cosmological parameters for predicting whether a cosmology leads to significant difference in M31’s velocities between realistic and traditional pairs. The left panel shows the scatter matrix for $v_{r}$, and the right panel for $v_{t}$.
@@ 6 Conclusion
In this work, we explored the barycenter velocity ($\mathbf{v}_{b}$) as a new kinematic constraint for defining Local Group (LG) analogues. This quantity has very low observational uncertainty but was not included in previous studies due to the need for sufficiently large simulation volumes to achieve numerical convergence in peculiar velocities (Forero-Romero & Sierra-Porta, 2022). Using simulations from the AbacusSummit project, with a box size of 2 Gpc $h^{-1}$, we have shown that including this constraint significantly influences the predicted velocity distributions of M31. We demonstrated this effect by comparing LG analogues selected using our proposed constraints (realistic pairs) with those selected using traditional constraints (separation distance less than 1 Mpc $h^{-1}$ and negative radial velocity). Our results show that in the realistic pairs, M31 tends to have a smaller radial velocity $v_{r}$ and larger tangential velocity $v_{t}$ compared to the traditional pairs. We quantified this effect using the difference in means and assessed its significance through two-sample Kolmogorov-Smirnov tests.
This effect is consistent across cosmological models and is particularly robust for $v_{r}$. In our analysis of 93 different cosmologies, we found significant differences in the radial velocity distribution for 86 models, while the tangential velocity distribution showed significant differences in 11 cosmologies. The mean differences between realistic and traditional pairs were 5% for $v_{r}$ and 1.4% for $v_{t}$, averaged across cosmologies.
We also found that $\sigma_{8\text{cb}}$ and $\Omega_{cdm}$ are the cosmological parameter that most strongly influences the significance of differences in radial and tangential velocity distributions respectively. However, further work is needed to understand the influence of cosmological parameters on LG kinematics. Similarly, while we have demonstrated statistical differences in velocity distributions, a deeper exploration of the physical mechanisms explaining why barycenter kinematics correlate with orbital properties remains for future work.
In summary, our results show that the kinematic constraints proposed in this work significantly influence the radial and tangential velocities of M31, with effects that are robust across cosmological models. This suggests that these new kinematic constraints should be considered when defining LG analogues in future studies.

### v3

@@ Abstract
We present a new way to identify systems similar to our Local Group (LG) of galaxies in cosmological simulations. Our method uses as a new constraint the speed and direction of the LG’s center of mass, which we can measure accurately from cosmic microwave background data. When we apply these criteria to different cosmological simulations and compare the results with traditional selection methods, we find statistically significant differences. Our approach produces simulated galaxy pairs where the relative M31 velocity is less radial ($2\%$ to $7\%$ difference over the mean) and more tangential ($1\%$ to $3\%$ difference over the mean) than in cases that do not take into account the barycenter speed. The radial change pattern appears consistently across all cosmological models we test, showing that matching the observed barycenter velocity has a measurable effect when modeling and interpreting Local Group-like systems.
@@ 1 Introduction
We identified an important measurement that previous studies haven’t used when defining LG-like systems. This measurement is the LG’s barycenter velocity ($\mathbf{v}_{b}$), which we can determine precisely from cosmic microwave background (CMB) data (Planck Collaboration et al., 2020). We use this constraint in two ways. First, we consider its magnitude ($v_{b}$). Second, we examine the direction, measured as the angle ($\theta$) between the barycenter velocity vector and the position vector of M31 relative to the MW ($\mathbf{r}$), as shown in Figure 1. We express this angle more conveniently as $\mu=\cos\theta$, which comes from the dot product of the unit vectors for $\mathbf{v}_{b}$ and $\mathbf{r}$.
Figure 1: Diagram showing the alignment between key quantities used to select LG analogues. $\mathbf{v}_{b}$ represents the barycenter velocity of the LG and $\mathbf{r}$ shows the position of M31 relative to the MW.
@@
Our paper is organized as follows: Section 2 covers the observational constraints, Section 3 describes the simulations used and the different cosmological models. In Section 4, we explain how we select LG analogues using both traditional criteria and our new approach. Section 5 presents and discusses our findings, and Section 6 summarizes our conclusions.
@@
$v_{r,\text{M31}}$ $109.3\pm 4.4$ $\text{kms}^{-1}$ van der Marel et al. (2012)
@@
$\mu$ $0.88\pm 0.01$ adim. This work
@@ 3 Simulation
For this study, we use the AbacusSummit simulation suite (Maksimova et al., 2021; Garrison et al., 2021). We analyze the base simulation box with $6912^{3}$ particles in a 2 Gpc/$h$ volume. Because the simulation snapshot is at redshift $z=0.1$, we use the halos’ peculiar velocities to project their positions to present-day values.
@@
Figure 2: Left panel: cumulative distribution of $\mu$, the cosine of the angle between the barycenter velocity of the LG and the position vector of M31 relative to the MW. We also plot a uniform distribution from -1 to 1 for comparison Right panel: Cumulative distribution for the barycenter speed $v_{b}$ of the Local Group.
@@ 4 Selection of LG analogues
We use halos identified with the CompaSO Halo Finder (Hadzhiyska et al., 2021), with selection criteria similar to those in Forero-Romero & Sierra-Porta (2022). First, we selected halos with maximum circular velocities $V_{\text{max}}>200$ km s-1. From these, we identify isolated pairs as two halos, A and B, that are each other’s nearest neighbors. We also require that no third halo with a maximum circular velocity greater than halo B’s can be found within three times the pair’s separation distance. We define halo A as the one with the greater maximum circular velocity, which we designated as M31. This forms our main sample, typically containing about $8\times 10^{5}$ pairs per simulation box.
@@
  • Total pair mass between $0.5,5.0]\times 10^{12}$ M
@@
  • Barycenter speed greater than the observed value, including the uncertainty ($v_{b}>605$ km s-1)
  • Alignment value less than the observed value, including the uncertainty ($\mu<-0.87$)
@@
Nonetheless, we include for completeness a variation of the realistic subsample, hereafter referred to as centered, which replaces the one-sided bounds with a symmetric interval around the observed mean values:
  • 500 $^{-1}<v_{b}<740 kms-$1
  • $0.80<\mu<-0.96$
The traditional subsample typically contains about $2\times 10^{5}$ pairs per simulation box, while the realistic and centered sample contain about $10^{4}$ pairs.
@@ 4.1 Statistical Tests
Figure 3: Mean velocity components of M31 relative to the MW as a function of $v_{b}$ and $\mu$ in the base cosmology (c000) for the traditional sample. The top panel shows $v_{r}$ and the bottom panel shows $v_{t}$
@@ 5.1 Base cosmology
Figure 2 shows the cumulative distributions of $\mu$ and $v_{b}$ in the base cosmology (c000). These distributions illustrate the key features of our barycenter velocity constraints used to select LG analogues. The $\mu$ distribution of reveals that $\mathbf{v}_{b}$ and $\mathbf{r}$ in simulated traditional LG pairs tend to be either aligned or anti-aligned. We confirm this by comparing against a uniform distribution using a KS test, which shows statistically significant differences ($p<0.05$). Using the same statistical method, we also find that the distribution is not symmetric with a slight tendency towards alignment.
@@
Figure 3 shows the mean $v_{r}$ as a function of $v_{b}$ and $\mu$ for the traditional pairs in the base cosmology. The figure indicates that smaller radial velocities correspond to higher values of $v_{b}$ and more extreme values of $\mu$ i.e. more aligned or anti-aligned pairs. In contrast, no clear relationship is observed for the tangential velocity.
@@ 5.2 All cosmologies
Figure 4: Mean values of M31 velocity distributions in all cosmologies (top panel shows $v_{r}$ and the bottom panel $v_{t}$). The main y-axis corresponds to the mean values of the velocities for traditional and realistic pairs. The secondary y-axis shows the percentage difference in mean values of realistic pair distributions compared to the mean of traditional pair distributions in each cosmology. Dashed lines and shaded regions represent the average and standard deviation of the mean values of traditional pair distributions for all cosmologies
Figure 5: Statistical significance of the differences between realistic and traditional selection criteria for radial velocity $v_{r}$ (top panel) and tangential velocity $v_{t}$ (bottom panel). The $p$-values shown are from two-sample Kolmogorov-Smirnov tests across all cosmologies.
Table 2: Number of cosmologies where the difference between traditional and realistic or centered pairs was statistically significant for each combination of observables.
@@
Observable N. cosmo. (Realistic) N. cosmo. (Centered)
None 6 47
$v_{r}$ only 65 34
$v_{t}$ only 1 9
$v_{r}$ and $v_{t}$ 12 3
$M_{\text{LG}}$ only 0 0
$v_{r}$ and $M_{\text{LG}}$ 7 0
$v_{t}$ and $M_{\text{LG}}$ 0 0
All three 2 0
Total 93 93
Our barycenter velocity constraints consistently affect the distributions of both $v_{r}$ and $v_{t}$ compared to traditional constraints across different cosmological models.
Figure 4 shows how the mean values of tangential and radial velocities change across all cosmologies. Realistic pairs consistently show less negative $v_{r}$, with mean differences between -2% to -7%. On the other hand, $v_{t}$ tends to increase for realistic pairs. The effect in this case is weaker, with changes between -1% and 3%.
This stronger effect in $v_{r}$ is also reflected in the significance tests. Figure 5 shows the p-values of the KS tests between traditional and realistic pairs velocity distributions. Among the 93 cosmologies we study, the differences in radial velocity distribution are significant in 86 models, while tangential velocity distribution show significant differences in 15 cosmologies. The effect on total mass ($M_{\text{LG}}$) is significant in only 9 cosmologies. Table 2 summarizes for each cosmology, which combinations of observables display statistically significant differences between traditional and realistic or centered pairs.
Table 3: Percentage differences in the distributions of the M31 velocity components ($v_{r}$ and $v_{t}$) and the total mass of the Local Group ($M_{\text{LG}}$). For each observable and sample type, the table reports the mean difference relative to the traditional sample, computed (i) using only the cosmologies where the difference is statistically significant and (ii) using all 93 cosmologies. The last column gives the number of cosmologies where the ks test indicates a significant difference between the distributions.
{ruledtabular}
Observable Sample Mean difference I (%) Mean difference II (%) Significant cosmologies (out of 93)
$v_{r}$ Realistic 4.64-4.64 4.51-4.51 86
Centered 2.83-2.83 2.14-2.14 37
$v_{t}$ Realistic $1.63$ $0.96$ 15
Centered 0.75-0.75 0.33-0.33 12
$M_{\text{LG}}$ Realistic 0.83-0.83 0.31-0.31 9
Centered - 0.06-0.06 0
Figure 6: Cumulative distributions of M31 velocities relative to the MW in cosmology 004 and cosmology 130, where we found the biggest difference in the mean between traditional and realistic pairs. In cosmology c004, the difference in both $v_{r}$ (-7%) and $v_{t}$ (3%) was statistically significant. In cosmology 130 only the difference in $v_{t}$ (3%) was statistically significant.
We calculated the average shift in the observables across cosmologies with significant differences, and to avoid selection bias, we repeated the procedure across all cosmologies. The results shown in table 3 further show the mentioned trend in the measured shift of the observables. The change in $v_{r}$ seen in the realistic sample is the largest and most robust across cosmologies. The mean change of roughly 4.6% in the significant subset remains almost the same (4.5%) when averaged over all cosmologies. The centered sample follows the same trend but with smaller magnitudes. Tangential velocity behaves differently. The mean difference in significant cosmologies is small (around 1.6% for the realistic sample), and when averaged over all cosmologies the value drops below 1%. For the centered sample the average change is negative (which correspond smaller velocities), but the magnitude is below 1% for the both averages (across significant cosmologies and across all cosmologies). $M_{\text{LG}}$ has the weakest effect, with changes below 1 percent and significant differences in only 9 cosmologies for the realistic sample and none for the centered.
[…]
Although the barycenter-based selection produces mean shifts that are statistically significant given the large sample sizes, the absolute shifts ($\sim$ a few km $s^{-1}$) are small compared to the typical velocity dispersion of LG analogues (tens of km $s^{-1}$); thus the effect is a modest kinematic refinement rather than a major physical change.
@@
Figure 7: Relative importance of various cosmological parameters for predicting whether the difference in M31’s velocities between centered and traditional pairs is statistically significant. The left panel shows the parameter importance for $v_{r}$, and the right panel for $v_{t}$. Importance is calculated using a Random Forest algorithm.
@@
Figure 8: Scatter plots of the mean difference in M31’s velocities between centered and traditional pairs versus the most important cosmological parameters for predicting whether a cosmology leads to significant differences. The left panel corresponds to $v_{r}$, and the right panel for $v_{t}$. For each parameter, $\rho$ is the calculated Pearson correlation coefficient with respect to the mean difference in the velocity.
5.3 Cosmological parameters
To explore which cosmological parameters most strongly influence whether the differences in $v_{r}$ and $v_{t}$ between traditional and realistic pairs are statistically significant, we trained a Random Forest classifier using the KS-test outcome (significant vs. not significant) as the target. However, the realistic filter introduces a class imbalance. For $v_{r}$, in 86 out of 93 cosmologies we found significant differences and for $v_{t}$, only in 11. Under these conditions, the classifier is dominated by the majority class and cannot reliably learn the minority class. This is reflected on the performance metrics obtained: 0.48 macro F1 score and weighted F1 score of 0.89.
For this reason we also considered the centered filter which defines a broader range around the mean observed values of $v_{b}$ and $\mu$. This filter produces a more balanced distribution of significant and non-significant cosmologies. In this case the F1 macro score was 0.55 and the weighted F1 score was 0.58. For this sample, the mean importance obtained for each parameter with a stratified K-fold validation is shown in figure 7. The most important parameter for determining significance for both velocity components was $A_{s}$ (scalar perturbation amplitude).
We then computed Pearson correlation coefficients between the mean velocity differences and individual cosmological parameters. Under the realistic filter, $v_{r}$ shows no measurable correlation, while $v_{t}$ shows a moderate negative correlation ($\rho\approx-0.45$) with both $A_{s}$ and $\sigma_{8}$. Under the centered filter however, the mean difference in $v_{r}$ exhibits moderate positive correlations ($\rho\approx 0.3$), with both $A_{s}$ and $\sigma_{8}$, while the correlation with $v_{t}$ remains unchanged.
The appearance of $A_{s}$ and $\sigma_{8}$ among the most influential parameters in the Random Forest Analysis, and their moderate correlations under the centered filter, suggests that the amplitude of matter fluctuations may play a role in setting how strongly barycenter kinematics affect Local Group analogue velocities. A higher fluctuation amplitude deepens gravitational potentials and may influence how coherent the barycenter motion is relative to halo orbits. Exploring these mechanisms, however, requires a dedicated physical study and lies beyond the scope of this work. The results presented here should therefore be regarded as exploratory, illustrating statistical trends rather than establishing physical causation.
@@ 6 Conclusion
In this work, we explored the barycenter velocity ($\mathbf{v}_{b}$) as a new kinematic constraint for defining Local Group (LG) analogues. This quantity has very low observational uncertainty but was not included in previous studies due to the need for sufficiently large simulation volumes to achieve numerical convergence in peculiar velocities (Forero-Romero & Sierra-Porta, 2022). Using simulations from the AbacusSummit project, with a box size of 2 Gpc $h^{-1}$, we have shown that including this constraint introduces a small but systematic shift in the predicted velocity distributions of M31. We demonstrated this effect by comparing LG analogues selected using our proposed constraints (realistic pairs) with those selected using traditional constraints (separation distance less than 1 Mpc $h^{-1}$ and negative radial velocity). Our results show that in the realistic pairs, M31 tends to have a smaller radial velocity $v_{r}$ and larger tangential velocity $v_{t}$ compared to the traditional pairs. We quantified this effect using the difference in means and assessed its significance through two-sample Kolmogorov-Smirnov tests.
This effect is consistent across cosmological models and is particularly robust for $v_{r}$. In our analysis of 93 different cosmologies, we found significant differences in the radial velocity distribution for 86 models, while the tangential velocity distribution showed significant differences in 15 cosmologies. The mean differences between realistic and traditional pairs were -4.5% for $v_{r}$ (less radial) and 1% for $v_{t}$ (more tangential), averaged across all cosmologies.
We also found that $A_{s}$ and $\sigma_{8}$ are the cosmological parameter that most strongly influences the significance of differences in radial and tangential velocity distributions. However, further work is needed to understand the influence of cosmological parameters on LG kinematics. Similarly, while we have demonstrated statistical differences in velocity distributions, a deeper exploration of the physical mechanisms explaining why barycenter kinematics correlate with orbital properties remains for future work. Assessing if these results change our understanding of the Local Group properties and history –for instance in light of studies such as Chamberlain et al. (2022), Hartl & Strigari (2024),Benisty & Mota (2025)– is also left for future work.
In summary, our results show that the kinematic constraints proposed in this work consistently influence the radial and tangential velocities of M31, with effects that are robust across cosmological models. Although small in magnitude, these systematic effects highlight that the LG’s bulk motion encodes useful information about its internal dynamics. Incorporating these new kinematic constraints should therefore be considered when defining LG analogues in future studies.